The measure of two complementary angles are in ratio 1 to 3. What are the measures of the angles?
Found 2 solutions by jim_thompson5910, addingup:
Answer by jim_thompson5910(35256)

solve for x in x+3x = 90

then use the value of x to compute the two angles (one angle is x, the other is 3x)

Answer by addingup(3677)

Complementary angles are either of two angles whose sum is 90°. So, your 2 angles add up to 90°.
1x+3x= 90
4x= 90
x= 90/4= 22.5
Angle 1 = 22.5
Angle 1*3= 22.5*3= 67.5
Check: 22.5+67.5= 90 Correct.